OpenStudy (anonymous):

A man sold 2 chairs at 240 dollars each. One at a profit of 20% and the other at a loss of 20%. On the whole he, a. broke even b. lost 20 c. gained 20. Explain pls

OpenStudy (anonymous):

lost

OpenStudy (anonymous):

pls explain why he lost 20. step by step solutions.

OpenStudy (anonymous):

so the original cost of the chair that profited is x+0.2=240 x=200 the original cost of the chair that lost is x-0.2x=240 x=300 total original cost=500 total money after selling 2 chairs = 240*2=480 so he lost 20 dollars

OpenStudy (anonymous):

i don't have load and i don't have skype in this computer. anyway he lost 20 he sold 2 chairs both at 240. but both were not the original price. let 1st original price be x; he got a profit of 20% meaning 1.2x 1.2x=240, x=200 meaning the original price was 200 but he sold it for 240 in the end let 2nd original price be y; he got a loss of 20% meaning .8y .8y=240, y=300 meaning the original price was 300 but he sold it for 240 in the end since he sold both at 240 he earned 480 but if you add both original costs, 200+300, it means he spent 500 so he lost 20 :D

OpenStudy (anonymous):

MEDAAAL. HAHAHA kidding, kidding =))